OpenStudy (anonymous):

Which number equals √147 3√7 7√2 7√3 3√3

OpenStudy (anonymous):

\[7\sqrt{3}\]

OpenStudy (anonymous):

\[7\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{147}=\sqrt{49*3}=7\sqrt{?}\]

OpenStudy (anonymous):

square the number outside of the √ the multiply the # inside